A 51 g sample of N2 (Nitrogen) gas has a volume of 32 L at 33°C.
i. Calculate the pressure using the Ideal Gas Law.
ii. Calculate the corrected pressure using the Van der Waals Equation.
a=1.351 atm×L2 b=0.0387 L
mol2 mol
1
Expert's answer
2020-04-15T01:53:06-0400

T=33+273=306 K

Mr(N2) = 14*2=28

n=51/28=1.82 mol

i. Ideal Gas Law : pV=m(N2)/Mr(N2) *RT

pi=m(N2)*R*T/(Mr(N2) *Vi)=(51*8.314*306) / (28*32)=144.8 kPa

ii.  Van der Waals Equation: [ pr + a*(n/V)2](Vr - n*b)=nRT

Vi=Vr - n*b, so [ pr + a*(n/V)2]Vi=nRT

pr= (nRT)/(V-n*b) - a*(n/V)2 = (1.82*8.314*306)/(32) - 1.351*(1.82/32)2 = 144.69 kPa